OpenStudy (anonymous):

the edges of a cube increases at a rate of 3 cm/s. how fast is the volume changing when the lenght of each edge is 60 cm?

OpenStudy (ipwnbunnies):

\[V = s^3\] Correct? When doing related rates, you want to take the derivative with respect to time, t, of the function.

OpenStudy (ipwnbunnies):

Sorry, I have to go but, taking the derivative with respect to t.... \[\frac{dV}{dt} = 3s^2 \frac{ds}{dt}\] Where 's' is the edge length and 'ds/dt' is the rate of change of the edge length. dV/dt will be the rate of change of the volume You're welcome.